OpenStudy (anonymous):

Two mechanics worked on a car. The first mechanic worked for 20 hours, and the second mechanic worked for 15 hours. Together they charged a total of $2100 . What was the rate charged per hour by each mechanic if the sum of the two rates was $125 per hour?

OpenStudy (anonymous):

\[M_{1}+M_{2}=$2100\]

OpenStudy (anonymous):

\[20x+15y=$125\]

OpenStudy (anonymous):

should i use one variable

OpenStudy (campbell_st):

well you have 20x + 15y = 2100 total cost and x + y = 125 total of the hourly rates solve simultaneously

OpenStudy (campbell_st):

x is the hourly rate for mechanic 1 y is the hourlyr ate for mechanic 2

OpenStudy (anonymous):

so i had them backwards deep

OpenStudy (anonymous):

ok thanks

OpenStudy (campbell_st):

sort of... you should always try to keep it as 2 unknowns in 2 equations.. they can be solved by elimination or substitution... whichever you feel most comfortable with

OpenStudy (anonymous):

aye thanks